A projectile is launched from ground level with the initial velocity of 50.0 m/s . The time of the flight is measured to be 6.00 s. How far from the launch point does the projectile land? Neglect the air resistance and g = 10.0 m/s^2


(A) 120 m

(B) 150 m

(C) 240 m

(D) 300 m


1
Expert's answer
2021-04-16T07:23:15-0400

From the time of the flight find the angle above the horizontal:


"\frac t2=\frac{v\sin\theta}{g},\\\space\\
\theta=\arcsin\frac{gt}{2v}."

Find how far the projectile landed:


"R=tv\cos\theta=tv\cos\bigg[\arcsin\frac{gt}{2v}\bigg]=240\text{ m}."